# The value of hedging against energy storage uncertainties when designing energy parks

### v1

@@ Abstract
Energy storage is needed to match renewable generation to industrial loads in energy parks. However, the future performance of bulk storage technologies is currently highly uncertain. Due to the urgency of decarbonization targets, energy park projects must be designed and begun now. But, as uncertainty in storage performance reduces, a different technology than identified during initial design may turn out cheaper. Enabling flexibility so that design adaptations can be made as better information becomes available would lower the cost of decarbonizing industry. But having this flexibility is itself costly. This raises the question, “Is it worth it?”
@@ 1 Introduction
1.1 Background
Decarbonizing electricity generation is critical to achieving 2050 net-zero carbon emissions targets [1]. Currently 40% electricity in the UK is generated from carbon intensive sources [2], and electricity demand is expect to increase by a factor 3 to 4 as heating and transport energy usage are decarbonized via electrification [3]. As a result, around 100 GW of wind and 60 GW of solar generation need to be constructed by 2050 [3]. However, managing the variability of electricity generation from these renewable sources to maintain security of supply presents a significant challenge [4], and the required adaptation of the energy system will be costly. Grid-scale energy storage is expected to be the main mechanism for matching renewable electricity supply with demand. Co-locating bulk energy storage with renewable generation and industrial loads, referred to as an energy park, has been proposed as a way of both improving the profitability of grid-scale storage (reducing the cost of its services), and reducing the grid impact of variable renewable generation [5, 6].
Many different energy storage technologies have been considered for bulk energy arbitrage, such as Lead-acid batteries, Lithium-ion batteries, compressed-air storage, and redox flow batteries [7]. However the future performance of these technologies, i.e. how they perform when implemented in the grid from around 2030 onwards, is still highly uncertain. Across the literature, estimates of technical parameters (e.g. round-trip efficiency and self-discharge rate) and economic parameters (e.g. cost per energy capacity and service lifetime) vary over wide ranges [7, 8, 9, 10, 11]. For example, a literature review, [8], found estimates for Sodium-Sulphur high-temperature batteries ranging from 70–90% for round-trip efficiency, and 0–20% for self-discharge rate. An IRENA report, [7], estimated the annualized energy capacity cost of compressed-air storage to be between 2 and 355 cents/kWh/yr.
@@ 1.2 Storage technology performance uncertainty in energy park design
The impact of uncertainty on the operation of energy parks, and so the cost of arbitraging energy to match variable generation with demand, has been widely studied [12, 13, 14, 15]. Many previous articles have investigated the optimal sizing of generation and storage assets to maximize total profit in the presence of various uncertainties. For example, Kim et al. [16] consider the sizing of wind generation, battery storage, and electrolyzer capacity in a micro-grid with local demands for both electricity and hydrogen. A bi-level optimization method is used to minimize the average cost of meeting the local energy demands, accounting for uncertainty in wind generation and demand patterns on operating costs, and uncertainty in the per capacity costs of the assets. Gabrielli et al. [17] develop a method for sizing a multi-energy system with solar generation, battery storage, heat pumps, and an electrolyzer, so that the design is robust to uncertainty in weather conditions and the electricity and heating demands which must be met. Bakke at al. [18] study the optimal level and timing of investment in stand-alone battery storage for providing energy arbitrage and ancillary grid services, considering uncertainty in both electricity prices (which determine operational revenue) and evolution in the price per capacity of battery storage.
Uncertainty in renewable generation [16, 17, 19, 20, 21, 22, 23, 24, 25], energy demands [16, 17, 21, 22, 23, 24, 25, 26], and the price of electricity [18, 21, 22, 23, 24] and carbon [19] are commonly considered during energy park design. However, few studies account for uncertainties related to storage technologies. Some have included uncertainty in the cost of energy storage [16, 18, 27, 22], however uncertainty in the technical performance of storage technologies, such as round-trip efficiency or self-discharge rate, is rarely considered [27, 22].
No existing works have investigated the impact that uncertainty reduction has on the process of energy park design, as has recently been studied in the context of building energy systems [28, 29]. Additionally, in the broader energy systems literature, risk aversion is known to be an important feature of decision making [30, 31, 32], but the impact it has on the benefit of uncertainty reduction is yet to be studied.
@@ 1.3 Storage technology selection in energy park design
While the majority of energy park design studies consider only a single storage option, a few do compare multiple storage technologies to identify the option best suited to the energy system considered [33, 34, 35]. Zhang et al. [33] compare the use of hydrogen and Lithium-ion batteries to manage the variability of wind generation and reduce curtailment. [34] & [35] compare different battery storage technologies for supporting distributed energy systems with local renewable generation. They find that Lead-acid and Sodium-Sulphur batteries provide the lowest total operating cost for their respective energy systems. However, none of these studies account for uncertainty in the characteristics of the storage technologies during the design optimizations which are then compared. A comparison of battery technologies which does account for performance uncertainties is performed in the context of building energy systems in [27].
The main limitation of all of these existing studies is that they consider the energy park design process, including both technology selection and sizing, as a static, single-point decision, where the entire design is determined up-front. This does not reflect the design of practical energy systems, where several stages of engineering design are performed progressively with increasing detail, and greater information regarding available components (which for instance may be provided during contracting).
As uncertainty in the performance of the available storage technology options reduces, the best technology choice (that with, for example, the lowest average cost under the remaining uncertainty) may change. As a result, retaining optionality over storage technologies through the design process is valuable, as it may be possible to reduce the overall cost of the energy system by selecting a better performing technology at a later stage. However, keeping this optionality and gathering improved information about the performance of each storage technology is costly. This raises the question of whether the benefits of retaining storage optionality are worth the cost.
@@ 1.4 Research objectives & novel contributions
This study uses the Value of Information analysis (VoI) framework to numerically answer the following questions, “How valuable is retaining optionality in storage technology choice for the design of an energy park as uncertainty reduces? Is the benefit to design worth the cost of obtaining improved estimates of storage performance?” It investigates this in the context of sizing wind, solar, and storage capacities for an illustrative energy park system, modelled off a real-world green hydrogen plant proposal in the Port of Rotterdam. The main objectives of this study are:
@@
    Quantify the benefit of adapting the system design (sizings) after uncertainty in storage performance has been reduced
@@
This work is the first to investigate the impact of reducing uncertainty in storage technology performance on energy system design, and quantify the benefit of retaining optionality in storage selection as uncertainty reduces. It demonstrates the value to energy system developers of providing the opportunity to adapt system design as improved information about the performance of energy storage (with reduced uncertainty) becomes available, which has important implications for industry design practice.
The remainder of this work is structured as follows. Section 2 outlines the Value of Information analysis framework used to study design under uncertainty reduction. Section 3 describes the illustrative energy park studied, including the probabilistic models of renewable generation, industrial load, and energy storage technology performance uncertainty used, and the Stochastic Programming model used to perform system design. Section 4 presents the results of the numerical experiments and discusses their importance for informing energy park design practices. An initial experiment quantifies the value of being able to adjust the energy park design, and choice of energy storage technology, after uncertainty in storage technology performance has been reduced by R&D. However, as no information is available in the literature regarding how much R&D might reduce uncertainty, i.e. how closely the performance of a demonstrator system matches a large-scale version, a sensitivity analysis over the level of uncertainty reduction is performed. The value of optionality is then investigated in two further cases where two energy storage technologies are used in the park, and where risk aversion is incorporated into the design process. Finally conclusions are drawn in Section 5.
@@ 2.1 Bayesian Decision Analysis
The actor, who is assumed to be risk neutral, aims to maximise the expected utility they receive from the ‘decision action’ they select. Costs are defined as negative utilities. The actor may choose to do this without taking any measurement. The resulting stochastic optimisation is termed the Prior Decision Problem,
@@
Alternatively, the actor can initially take a ‘measurement action’ which reduces uncertainty in the problem, and improves the following choice of ‘decision action’. In this case, the optimization of expected utility is performed over both measurement and decision actions, and is termed the Pre-Posterior Decision Problem,
@@ 3 Experimental setup
3.1 Illustrative energy park system
[…]
@@ 3.2 System design under uncertainty
A key design challenge is determining the sizing of generation and storage, and the energy storage technology, that maximize the profitability of the energy park. It is assumed that offshore wind and solar generation (via direct wire connection) can be installed to support the site, and that energy can be traded with the grid. Fig. 2 illustrates the energy flows within the system model. Limits are imposed on grid connection capacity to reflect network restrictions, and on the maximum solar generation capacity to reflect land scarcity. Additionally, a capital budget constraint is imposed. Optimization-based design is used to select the energy storage technology and generation & storage sizings that minimize the expected cost of providing the plant load. A carbon price is included to reflect the incentives for low-carbon hydrogen production. Uncertainties in the industrial load, patterns of wind and solar generation, and cost and technical performance of the different storage technologies are accounted for during design.
@@
This decision model raises a question, “Even though only one storage technology is ultimately used in the energy park, is it worthwhile performing procurement and R&D for all technologies, so that at the final design stage a possibly better choice of storage technology can be made once uncertainty has been reduced?”, i.e. “Is retaining optionality in storage technology choice until the final design stage worth the cost?”.
3.3 Probabilistic models of uncertainties
3.3.1 Renewable generation & industrial load
Historic data was used to construct probabilistic models of uncertain wind and solar generation. In both cases, generation potential per installed capacity data was collected for 2010-2019. Uncertainty in renewable generation was modelled by randomly sampling a year from that range and using the corresponding generation data. For wind, data was collected from the renewables.ninja model [39, 40] at the location of a wind farm development zone near Rotterdam (IJmuiden Ver), using a power model from a typical offshore wind turbine. For solar, data was collected from the EU’s PVGIS model [41], using default PV module settings. Fig. 4 plots this data from periods in both summer and winter for three example years.
@@ 3.3.1 Renewable generation & industrial load
[…]
The probabilistic models of the renewable generation and industrial load time series are summarized in Table 1. Values from 2023 for the Netherlands were used for the price and carbon intensity of grid electricity [^1] Further information on data collection and processing is available online.. Uncertainties in price and carbon intensity were not modelled due to a lack of multiple years of representative data.
@@
[…]
@@
3.3.2 Storage technology performance & measurement
@@
All uncertain parameters are taken to be independent, with mean ($\mu$) and standard deviation ($\sigma$) values derived separately for each parameter from literature estimates [7, 8, 11, 9].
@@
The uncertainty reduction factor, $r$, models the confidence that the energy park designer has in the demonstrator, i.e. how closely they believe the performance of the demonstrator reflects that of a large-scale system.

### v4

@@ Abstract
Energy storage is needed to match renewable generation to industrial loads in energy parks. However, the future performance of bulk storage technologies is currently highly uncertain. Due to the urgency of decarbonization targets, energy park projects must be designed and begun now. But, as uncertainty in storage performance reduces, a different technology than identified during initial design may turn out cheaper. Enabling flexibility so that designs can be updated as better information becomes available would lower the cost of decarbonizing industry. But having this flexibility is itself costly. This raises the question, “Is it worth it?”
@@ 1 Introduction
Decarbonizing electricity generation is critical to achieving 2050 net-zero carbon emissions targets [9]. Currently 40% of electricity in the UK is generated from carbon intensive sources [14], and electricity demand is expect to increase by a factor 3 to 4 as heating and transport energy usage are decarbonized via electrification [32]. As a result, around 100 GW of wind and 60 GW of solar generation need to be constructed by 2050 [32]. However, managing the variability of electricity generation from these renewable sources to maintain security of supply presents a significant challenge [34], and the required adaptation of the energy system will be costly. Grid-scale energy storage is expected to be the main mechanism for matching renewable electricity supply with demand. Co-locating bulk energy storage with renewable generation and industrial loads, referred to as an energy park, has been proposed as a way of both improving the profitability of grid-scale storage (reducing the cost of its services), and reducing the grid impact of variable renewable generation [7].
The design of energy parks, including the selection and sizing of renewable generation and energy storage technologies, has been extensively studied in the literature. For example, Arévalo & Jurado [1] investigates the combination of energy storage technology and control scheme that provides the lowest net present cost for supporting an autonomous grid with solar, wind, hydro, and diesel generation. Sizing optimization is performed for each system configuration, and it is found that due to the different performance characteristics of the storage technologies, they are best suited to supporting substantially different generation mixes. This indicates that the technical performance of energy storage has a significant impact on how the energy system should be sized. Hu et al. [20] considers a district energy system with local renewable generation providing both electricity and heat. It demonstrates that optimizing the sizing and location of battery storage, hydrogen storage, and CHP units in the system can reduce the net load fluctuation by 25%. Similar benefits of using energy storage to manage variability in both renewable generation and load are shown by Phu et al. [36], which studies an energy park producing green hydrogen from biomass using electricity from wind and solar generation. Optimizing battery and hydrogen energy storage within the system is shown to substantially reduce the operating cost, carbon emissions, and grid reliance of the energy park. Further this optimized configuration is demonstrated to be robust against uncertainties in hydrogen demand, grid electricity price, and renewable generation.
Many different energy storage technologies have been considered for supporting energy parks, such as Lead-acid batteries, Lithium-ion batteries, compressed-air storage, and redox flow batteries [21]. However the future performance of these technologies, i.e. how they will perform when implemented in the grid from around 2030 onwards, is still highly uncertain. Across the literature, estimates of technical parameters (e.g. round-trip efficiency and self-discharge rate) and economic parameters (e.g. cost per energy capacity and service lifetime) vary over wide ranges [33]. For example, a literature review, [23], found estimates for Sodium-Sulphur high-temperature batteries ranging from 70–90% for round-trip efficiency, and 0–20% for self-discharge rate. [35] found values for the capital cost of Li-ion batteries between 150 and 600 €/kWh, and operational costs between 1 and 5%. And an IRENA report, [21], estimated the annualized energy capacity cost of compressed-air storage to be between 2 and 355 cents/kWh/yr.
@@ 1.1 Storage technology performance uncertainty in energy park design
The impact of uncertainty on the operation of energy parks, and so the cost of arbitraging energy to match variable generation with demand, has been widely studied. For instance, [24] shows that accounting for load forecast uncertainty in control can reduce operating costs by 10%, and [22] develops a method for managing the risk introduced by uncertainty in load and solar generation during operation. Many previous articles have investigated the optimal sizing of generation and storage assets to maximize total profit in the presence of various uncertainties. For example, Kim et al. [25] consider the sizing of wind generation, battery storage, and electrolyser capacity in a micro-grid with local demands for both electricity and hydrogen. A bi-level optimization method is used to minimize the average cost of meeting the local energy demands, accounting for uncertainty in wind generation and demand patterns on operating costs, and uncertainty in the per capacity costs of the assets. Gabrielli et al. [16] develop a method for sizing a multi-energy system with solar generation, battery storage, heat pumps, and an electrolyser, so that the design is robust to uncertainty in weather conditions and the electricity and heating demands which must be met. Bakke at al. [3] study the optimal level and timing of investment in stand-alone battery storage for providing energy arbitrage and ancillary grid services, considering uncertainty in both electricity prices (which determine operational revenue) and evolution in the price per capacity of battery storage.
Uncertainty in renewable generation [16], energy demands [29], and the price of grid electricity [46] and carbon [47] are commonly considered during energy park design. However, few studies account for uncertainties related to storage technologies. Some have included uncertainty in the cost of energy storage, e.g. [25] & [3]. However only two studies could be found that consider uncertainty in the technical performance of storage technologies, [10] which accounted for uncertain cycle life and self-discharge rate, and [6] which considered uncertain round-trip efficiency.
No existing works have investigated the impact that uncertainty reduction has on the process of energy park design, as has recently been studied in the context of building operation [27] and district energy system design [26]. Additionally, in the broader energy systems literature, risk aversion is known to be an important feature of decision making [38]. Its effect on the selection and sizing of energy infrastructure in parks has been investigated in studies such as [31] and [45] respectively. However, the impact risk aversion has on the benefit of uncertainty reduction is yet to be studied.
@@ 1.2 Storage technology selection in energy park design
While the majority of energy park design studies consider only a single storage option, a few do compare multiple storage technologies to identify the option best suited to the energy system. Zhang et al. [49] compares the use of hydrogen storage and Lithium-ion batteries to manage the variability of wind generation and reduce curtailment. [28] & [30] compare different battery storage technologies for supporting distributed energy systems with local renewable generation. They find that Lead-acid and Sodium-Sulphur batteries provide the lowest total operating cost for their respective energy systems. However, none of these studies account for uncertainty in the characteristics of the storage technologies during the design optimizations which are then compared. A comparison of battery technologies which does account for performance uncertainties is performed in the context of building energy systems in [6].
The main limitation of all these existing studies is that they consider the energy park design process, including both technology selection and sizing, as a static, single-point decision, where the entire design is determined up-front. This does not reflect the design process of practical energy systems, where several stages of engineering design are performed progressively with increasing detail, and greater information regarding available components (which for instance may be provided during contracting).
As uncertainty in the performance of the available storage technology options reduces, the best technology choice (that with, for example, the lowest average cost under the remaining uncertainty) may change. As a result, retaining optionality over which storage technology is ultimately used through the design process could be valuable, as it may be possible to reduce the overall cost of the energy system by selecting a better performing technology at a later stage. However, keeping this optionality and gathering improved information about the performance of each storage technology is costly. This raises the question of whether the benefits of retaining storage optionality are worth the cost.
@@ 1.3 Research objectives & novel contributions
This study uses the Value of Information analysis (VoI) framework to numerically answer the following questions, “How valuable is retaining optionality in storage technology choice for the design of an energy park as uncertainty reduces? Is this benefit to design worth the cost of obtaining improved estimates of storage performance and retaining the option to change storage technology?”. It investigates this in the context of sizing wind, solar, and storage capacities for an illustrative energy park system, modelled off a real-world green hydrogen plant proposal in the Port of Rotterdam. The main objectives of this study are:
@@
    Quantify the benefit of updating the system design (sizings) after uncertainty in storage performance has been reduced
@@
This work is the first to investigate the impact of reducing uncertainty in storage technology performance on energy system design, and quantify the benefit of retaining optionality in storage technology selection as uncertainty reduces. It demonstrates the value to energy system developers of providing the opportunity to update system designs as improved information about the performance of energy storage (with reduced uncertainty) becomes available, which has important implications for industry design practice.
The remainder of this work is structured as follows. Section 2 outlines the Value of Information analysis framework used to study design under uncertainty reduction. Section 3 describes the illustrative energy park studied, including the probabilistic models of renewable generation, industrial load, and energy storage technology performance uncertainty used, and the Stochastic Programming model used to perform system design. Section 4 presents the results of the numerical experiments and discusses their importance for informing energy park design practices. An initial experiment quantifies the value of being able to update the energy park design, and change choice of energy storage technology, after uncertainty in storage technology performance has been reduced by R&D. However, as no information is available in the literature regarding how much R&D might reduce uncertainty, i.e. how closely the performance of a demonstrator system matches a large-scale version, a sensitivity analysis over the level of uncertainty reduction is performed. The Value of Optionality is then investigated in two further cases where two energy storage technologies are used in the park, and where risk aversion is incorporated into the design process. Finally conclusions are drawn in Section 5.
@@ 2 Methodology
Value of Information analysis, originally proposed by Raiffa [39] and Howard [19] in the 1960s, is a framework based on Bayesian Decision Analysis and Expected Utility Theory [42] for quantifying the improvement in decision making provided by uncertainty reduction. This section briefly outlines the methodology, and shows how it can be extended to study the benefit of decision optionality as uncertainty reduces.
@@ 2.1 Bayesian Decision Analysis
The actor, who is assumed to be risk neutral, aims to maximise the expected utility they receive from the ‘decision action’ $a$ they select. Costs are defined as negative utilities. The actor may choose to do this without taking any measurement. The resulting stochastic optimisation is termed the Prior Decision Problem,
@@
Alternatively, the actor can initially take a ‘measurement action’ $e$ which reduces uncertainty in the problem, and improves their subsequent choice of ‘decision action’. In this case, the optimization of expected utility is performed over both measurement and decision actions, and is termed the Pre-Posterior Decision Problem,
@@ 3 Experimental setup
The value of retaining optionality in energy storage technology choice during design is investigated for an illustrative energy park system. This system is modelled on a real-world energy park proposal. Due to its importance for decarbonizing transportation, heavy industry, and chemical manufacturing, there has been substantial commercial interest in the production of low-carbon hydrogen at scale. A proposal to build a $250\text{\,}\mathrm{M}\mathrm{W}$ hydrogen electrolyser facility in the Port of Rotterdam to decarbonize industry and transportation in the area has received early-stage support from the Dutch government [5]. Co-locating renewable generation and energy storage with this industrial hydrogen plant in an energy park would reduce the total cost of operating the system, and its impact on the electricity grid.
@@ 3.1 System design under uncertainty
A key design challenge is determining the sizing of generation and storage, and the energy storage technology, that maximize the profitability of the energy park. It is assumed that offshore wind and solar generation (via direct wire connection) can be installed to support the site, and that energy can be traded with the grid. Fig. 2 illustrates the energy flows within the system model. Limits are imposed on grid connection capacity to reflect network restrictions, and on the maximum solar generation capacity to reflect land scarcity. Additionally, a capital budget constraint is imposed. Optimization-based design is used to select the energy storage technology and generation & storage sizings that minimize the expected cost of providing the industrial load. A carbon price is included to reflect the incentives for low-carbon hydrogen production. Uncertainties in the industrial load, patterns of wind and solar generation, and cost and technical performance of the different storage technologies are accounted for during design.
@@
This decision model raises a question, “Even though only one storage technology is ultimately used in the energy park, is it worthwhile performing procurement and R&D for all available technologies, so that at the final design stage, a possibly better choice of storage technology can be made once uncertainty has been reduced?”, i.e. “Is retaining optionality in storage technology choice until the final design stage worth the cost?”.
3.2 Probabilistic models of renewable generation & industrial load
Historic data is used to construct probabilistic models of uncertain wind and solar generation. In both cases, generation potential per installed capacity data was gathered for 2010-2019. Uncertainty in renewable generation is modelled by randomly sampling a year from that range and using the corresponding generation data. For wind, data was collected from the renewables.ninja model [43] at the location of a wind farm development zone near Rotterdam (IJmuiden Ver), using a power model from a typical offshore wind turbine. For solar, data was collected from the EU’s PVGIS model [8], using default PV module settings. Fig. 4 plots this data from periods in both summer and winter for three example years.
@@ 3.2 Probabilistic models of renewable generation & industrial load
Uncertainty in the demand for hydrogen in the local market during operation leads to uncertainty in the final sizing of the electrolyser. The conversion efficiency of the electrolyser units also has some uncertainty. So, at the time the supporting generation and storage infrastructure is designed, the electrical load that will be required by the hydrogen plant is uncertain. The industrial load is modelled as being constant during operation, as electrolysers are typically run at high capacity factors to provide a competitive hydrogen price [15], but with an unknown/uncertain level. A truncated Gaussian distribution is used for the load, with mean set to the proposed electrolyser capacity of $250\text{\,}\mathrm{M}\mathrm{W}$, and standard deviation $\sigma$ taken to be $25\text{\,}\mathrm{M}\mathrm{W}$ (10%), truncated at $\pm 2\sigma$.
The probabilistic models of the renewable generation and industrial load time series are summarized in Table 1. Grid electricity price and carbon intensity data for the Netherlands from 2023 is used as a base case [^1] Further information on data collection and processing is available online.. Uncertainties in price and carbon intensity are not modelled, as due to global energy market events, few years of representative data are available. However 2024 data is used to perform a sensitivity analysis to determine if changes in grid price and carbon impact the conclusions derived.
@@
Load Truncated Gaussian $\mu=$250\text{\,}\mathrm{M}\mathrm{W}$\$ $\sigma=$25\text{\,}\mathrm{M}\mathrm{W}$\$ $\text{cut-off}=\pm 2\sigma$ Truncation used to limit unrealistic extreme load values
@@
3.3 Probabilistic models of storage technology performance & measurement
@@
All uncertain parameters are taken to be independent, with mean ($\mu$) and standard deviation ($\sigma$) values derived separately for each parameter from literature estimates.
@@
The uncertainty reduction factor, $r$, models the confidence that the energy park designer has in the demonstrator, i.e. how closely they believe the performance of the demonstrator reflects that of a large-scale system. Determining a precise numerical value for this parameter is challenging, as there is very little data available regarding how the performance of demonstrator systems compares to large-scale energy storage plants, particularly in the case of newer technologies such as CAES and VRFB where few large-scale projects have been completed. An initial estimate of $r=0.25$ is used for the experiments, representing a case where the designer has good confidence in the demonstrator, but retains a reasonable fraction of the initial uncertainty. To overcome this limitation, a sensitivity analysis over the value of $r$ is performed. Note that because $r$ is a multiplicative factor, the more mature storage technologies with lower uncertainty in their prior distributions also have correspondingly lower uncertainties in the posteriors.